OpenStudy (anonymous):

sin(−t) = 8/9 Evaluate sin t.

OpenStudy (anonymous):

sin(-x) = -sin(x)

OpenStudy (mathmate):

Sinus is an odd function, so it satisfies the following property. Sin(t)=-sin(-t) You can use the property to find sin(x) when sin(-x)=8/9.

OpenStudy (anonymous):

no need to find t if all yo want is sin(t)

OpenStudy (anonymous):

??

OpenStudy (anonymous):

if \(\sin(t)=\frac{3}{5}\) then \(\sin(-t)=-\frac{3}{5}\)

OpenStudy (anonymous):

if \(\sin(-t)=\frac{8}{9}\) then \(\sin(t)=-\frac{8}{9}\) that is all there are not two answers

OpenStudy (anonymous):

exactly...

OpenStudy (anonymous):

@satellite73 @pgpilot326 Sorry was chatting with prime didn't look carefully

OpenStudy (anonymous):

no worries just so @lesliemariahp gets it

OpenStudy (anonymous):

My homeworks are online and for some reason it marks this answer wrong. :(

OpenStudy (anonymous):

sin t = -8/9 is wrong?

OpenStudy (anonymous):

yup! @pgpilot326

OpenStudy (anonymous):

something's goofed in the problem statement or in the program online. sorry

OpenStudy (mathmate):

Does it require you to express it in decimal to so many places? Read all the instructions.

OpenStudy (anonymous):

Thank you guys regardless.

OpenStudy (anonymous):

@lesliemariahp did the cos(217) work last time?

OpenStudy (anonymous):

nope :(

OpenStudy (anonymous):

Weird...what format solutions did it accept before?

OpenStudy (anonymous):

it didn't accept any. Not decimal or fraction. Ugh, so frustrating. Usually is takes both.

OpenStudy (mathmate):

Check the instructions.

OpenStudy (anonymous):

screenshot?

OpenStudy (anonymous):

Here try this..